OpenStudy (anonymous):

solve: sqrt{7-2x}=x-2

OpenStudy (anonymous):

Answer is 3

OpenStudy (anonymous):

To get 3

OpenStudy (anonymous):

what is the work for it?

OpenStudy (anonymous):

You take the sqrt of both sides

OpenStudy (anonymous):

You will be left with 7-2x= x^2 -4

OpenStudy (anonymous):

Next combine like terms

OpenStudy (anonymous):

x^2-2x-11

OpenStudy (anonymous):

x^2+2x-11

OpenStudy (anonymous):

I made a mistake somewhere

OpenStudy (anonymous):

Let me check

OpenStudy (anonymous):

I should really do these things on paper

OpenStudy (dumbcow):

@Microrobot , please type more than 1 step in each post and you made a mistake squaring right side

OpenStudy (anonymous):

I did the same thing before too

OpenStudy (tkhunny):

@Microrobot You don't seem to be getting the idea. DO NOT just dump answers. The fact that the OP had to ask you to show your work exposes the pathology. The Code of conduct is important.

OpenStudy (anonymous):

Thank you dumbcow

OpenStudy (anonymous):

I am explaining the answer @tkhunny

OpenStudy (anonymous):

is it positive 4 instead of negative

OpenStudy (anonymous):

I made a mistake somewhere

OpenStudy (anonymous):

So im doing it on paper to prevent errors

OpenStudy (anonymous):

and then reexplain

OpenStudy (tkhunny):

\(\sqrt{7-2x} = x-2\) First, observe that \(7 - 2x \ge 0\) or \(2x \le 7\) or \(x \le 7/2\) WHY? Second, observe that \(x - 2 \ge 0\) or \(x \ge 2\) WHY?

OpenStudy (tkhunny):

@Microrobot Your first two posts were just answers. EIGHT of your other posts are just chatter. Just take a little more time and be sure of what you are doing. So far, you're quite a bit more on the confusing side of things.

OpenStudy (anonymous):

Ok you can take this one

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Sorry about that, thank you for informing me

OpenStudy (dumbcow):

@sydpendleton ,u still need any help?

OpenStudy (anonymous):

yes please

OpenStudy (dumbcow):

ok @tkhunny showed you the domain restrictions which tells you the solution must be in interval 2<x<7/2 @Microrobot , was on right track...square both sides, then solve the quadratic \[(\sqrt{7-2x})^{2} = (x-2)^{2}\] \[7-2x = x^{2} -4x+4\] \[x^{2} -2x-3 = 0\] \[(x-3)(x+1) = 0\] only 1 solution in domain....x=3

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

syd badge dumbcow

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Select Best Response for dumbcow

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I did :)

OpenStudy (anonymous):

Thanks!